Vern has investments totaling $11,000 in two accounts. Once account is a savings account paying 6% interest and the other is a bond paying 9%. If the annual interest from the two investments was $855: a. How much did he have invested at the 6% rate?
A) $4,500 B) $6,500 C) $6,600 D) $9,900
b. How much did he have invested at the 9% rate?
A) $4,500 B) $6,500 C) $6,600 D) $9,900
c. How much interest did the 6% account earn?
A) $270 B) $585 C) $660 D) $990
d. How much interest did the 9% account earn?
A) $270 B) $585 C) $660 D) $990


Answer by checkley71(8403)

.06X+.09(11000-X)=855 OR .06X+990-.09X=855 OR -.03X=855-990 OR .03X=135 OR
X=135/.03 OR X=4500 & 11000-4500=6500. 4500 INVESTED @ 6% & 6500 INVESTED @9%
.06*4500=270 INTEREST FOR THE 6% INVESTMENT
.09*6500=585 FOR THE 9% INVESTMENT